OpenStudy (broskishelleh):

Let f(x) = 7x - 13. Find f^-1(x).

OpenStudy (broskishelleh):

13x-7 7/x+13 x+13/7 1/7x-13

OpenStudy (jdoe0001):

well.. to get the inverse "relation", we simply first swap about the variables..... thus \(\bf f(x) ={\color{brown}{ y}}= 7{\color{blue}{ x}} - 13\qquad inverse\implies {\color{blue}{ x}}= 7{\color{brown}{ y}} - 13\) and solve that for "y"

OpenStudy (solomonzelman):

shouldn't it be +13 ?

OpenStudy (broskishelleh):

@jdoe0001 @SolomonZelman Isn't the answer y = 1/7x +13/7

OpenStudy (broskishelleh):

@Frostbite

OpenStudy (jdoe0001):

hmmm what did you get after solving for "y"?

OpenStudy (broskishelleh):

y = 1/7x +13/7

OpenStudy (broskishelleh):

Weird answer, huh?

OpenStudy (solomonzelman):

I just thought that + would be the inverse action to -.

OpenStudy (jdoe0001):

hmmm you may want to recheck your simplification

OpenStudy (broskishelleh):

Alright one sec

OpenStudy (broskishelleh):

I have come to another dead end as I got: y=7x−13

OpenStudy (jdoe0001):

x = 7y - 13 try "adding" 13 to both sides then divide both by 7

OpenStudy (solomonzelman):

I see what you did, jdoe, I misinterpreted you.. of course you are right (as always)

OpenStudy (broskishelleh):

So lets see: x + 13 = 7y - 13 + 13 13x = 7y --------- 7 7 13x --- = y 7

OpenStudy (broskishelleh):

13x/7

OpenStudy (jdoe0001):

hmm x+13 = 13x?